OpenStudy (anonymous):

Solve the system by substitution. 2x-y+z=-4 z=5 -2x+3y-z+-10 Possible answers: 1. -8,7,5 2. -8,-7,5 3. 8,-7,5 4. -8,-7,-5 I'm so confused I don't even know where to start please help!

OpenStudy (shamil98):

z=5 z is already given, so just solve the equation as if it were this 2x- y + 5 = -4 -2x +3y - 5 = -10 add the equations 2y = -14 y = -7 2x +7 + 5 = -4 2x + 12 = -4 2x = -16 x = -8

OpenStudy (shamil98):

Do you understand? or should I explain it a bit more?...

OpenStudy (anonymous):

Yeah I'm really bad at math I'm still a bit confused @shamil98

OpenStudy (shamil98):

Okay. so you have 3 equations 2x-y+z=-4 -2x+3y-z = -10 z=5 you see the z variables in the other two equations? you just input the 5 since z = 5. This then becomes 2x - y + 5 = -4 -2x +3y -(5)= -10 So far so good?...

OpenStudy (anonymous):

Okay I understand that part

OpenStudy (shamil98):

2x -2x = 0 so that means you can solve for y so just add these equations 2x - y + 5 = -4 -2x +3y -(5)= -10 ----------------- 0 +2y + 0 = -14 2y = -14 2y/2 = -14/2 y = -7

OpenStudy (shamil98):

You know have two variables done, just x to go. so you input those values in one of the original equation. 2x-y+z=-4 2x -(-7) +(5) = -4

OpenStudy (anonymous):

then add them or?

OpenStudy (shamil98):

2x + 7 + 5 = -4 2x + 12 = -4 -12 -12 2x = - 16 2x/2 = -16/2 x = -8

OpenStudy (shamil98):

Yes, you add that and solve for x.

OpenStudy (anonymous):

So the second choice -8,-7,5?

OpenStudy (shamil98):

Yep.

OpenStudy (anonymous):

Thank you so much I actually understood most of what you said(: